Use the following financial information for questions 2-4 below: From the income statement: Depreciation expense Interest expense Income tax Net income $112,000 25,000 45,000 53,000 From the balance sheet: Current liabilities Long-term debt Deferred income taxes Total Liabilities Preferred stock Common stock Premium on common stock Retained earnings Total Stockholders Equity Total Liabilities &Stockholders Equity $80,000 400,000 15,000 495,000 75,000 157,000 185,000 245,000 2. What is the Times Interest Earned ratio? 3. What is the Debt/Assets (Debt) ratio? 4. What is the Debt/Equity ratio?
Answer #1
2)Calculation of time interest earned ratio:
Time Interest Earned Ratio= Earning before Interest and tax/ interest expense
Earning before interest and tax= Net income+tax+interest expense
                                                               =53000+45000+25000= $123000
Time Interest Earned Ratio= 123000/25000=4.92
3)Calcualtion of debt/ assets ratio:
Debt/assets ratio= Total liabilities/Total assets
Total assets= Total liabilities & stockholders equity=$1157000
Debt/assets ratio=495000/1157000=0.43
4)Calcualtion of debt/equity ratio:
Debt/equity ratio= Total liabilities/stockholder's equity
Debt/equity ratio=495000/662000=0.75